How much energy is released when 29.7g of methane are burned? ΔH = - 890.4 kJ
CH4+2O2 > CO2+2H2O

A 236 gram sample of an unknown substance at 89.00°C was added to 52 grams of water at 23.00°C in a perfectly insulated calorimeter. The final temperature of the unknown substance-water mixture was 39.50°C. Calculate the specific heat of the unknown substance using this data.
1
Expert's answer
2017-05-29T09:55:10-0400
Solution 1:
n (〖CH〗_4 )= 29.7/16=1.86 (mol)
Q=1.86×|-890.4|=1652.81 (kJ)

Solution 2:
Q=c_1 m_1 ∆T_1=c_x m_x ∆T_x⇒ c_x=(c_1 m_1 ∆T_1)/(m_x ∆T_x )
c_x=(4.18×52×(39.50-23.00))/(236×(89.00-39.50))=0.307 (J/gK)
Answer: 0.307 Jg-1K-1.